# Monocular Lane Detection Based on Deep Learning: A Survey

### v2

@@ IV-B BEV-based methods
where $F_{FV}$ denotes the FV feature, $F_{BEV}$ denotes the BEV feature which contains information of height, $x$, $y$, $z$ denote coordinates in 3D space, $M_{trans}$ denotes view transformation module, $u$, $v$ denote corresponding pixel coordinates in terms of $x$, $y$, $z$. $[\bm{R\quad T}]$ and $\bm{K}$ are camera extrinsics and intrinsics.
@@ IV-B1 IPM for View Transformation
3D-LaneNet [22] is the first method which uses deep learning to predict 3D lanes directly from monocular images. The network first predicts the camera pitch angle and height to generate a differentiable IPM, combining the original FV feature map to create the BEV feature map. In the BEV plane, lanes are represented as vertical line anchors. 3D-LaneNet demonstrates promising results in detecting 3D lanes from monocular images. Gen-LaneNet [23] directly uses 2D lane segmentation results as input for IPM, allowing for the utilization of extensive 2D lane data and enhancing the model’s generalization. In contrast to 3D-LaneNet, Gen-LaneNet offers more reliable supervision by using the camera pitch angle and height as GT. Li et al. [118] propose a new loss function based on Gen LaneNet to better extract the height information of 3D lanes from 2D lane representations. In the BEV plane, the above work uses vertical line anchors to model lanes. However, matching line anchors to ground truth is performed by measuring the distance at a predefined $y=Y_{ref}$, which may result in missed detections for short lanes. Therefore, 3D-LaneNet+ [129] avoids this issue using a bottom-up segmentation approach. Liu et al. [130] believe that the model can be independent of the ground truth camera pose provided by the benchmark. They designed a two-stage network based on Transformer, which first predicts camera pose, i.e., the required parameters for IPM, then extracts BEV features, and finally regresses polynomial coefficients. 3D-SplineNet [122] treats lanes as B-spline curves in the BEV plane.
@@ IV-C1 Combining Depth Estimation
Similar to depth-assisted monocular 3D object detection methods [134, 135, 136], SALAD [44] decouples 3D lane detection into 2D lane segmentation and dense depth estimation tasks. With the help of estimated depth, it projects 2D lane coordinates into 3D space. Due to the availability of depth information, this method is independent of camera extrinsic.
TABLE IV: Benchmark results of 2D lane detection methods on CULane [36]. For “Cross” category , only false positives are shown.
@@ IV-C2 Derictly Modeling 3D Lanes
The 3D object detection methods based on sparse BEV representations, such as DETR3D [137] and PETR [138], guide this approach. CurveFormer [119] constructs curve queries by modeling 3D lanes as 3D line anchors to provide explicit positional priors [139]. It designs a curve cross-attention mechanism to predict polynomial parameters for 3D lanes. Inspired by the line anchor feature pooling mechanism in LaneATT [80], Anchor3DLane [25] utilizes camera intrinsics and extrinsics to accurately project 3D line anchor points onto FV features. This facilitates anchor feature sampling, allowing the network to predict 3D coordinates and lane classification results based on the sampled anchor features. LATR [27] decomposes 3D line anchors into dynamically generated point-level and lane-level queries. It uses dynamic 3D ground position embeddings to interact with FV features, updating the lane query to bridge 3D space and 2D images. Dong et al. [124] model lanes as 3D Bézier curves and predict the curve control points via Transformer. Han et al. [140] expressed the 3D lane as a polynomial in 3D space. They design two Transformer structures to learn the 2D polynomial and height of the X-O-Z plane and project the resulting 3D lane onto FV for supervised alignment. PVALane [123] generates sparser 3D line anchors than Anchor3DLane by first predicting 2D lanes in FV. It also introduces a module to align sampled FV and BEV features for more accurate 3D lane detection.
@@ V Benchmark Results
This section reports the performance of representative lane detection methods on commonly used public datasets. For each reviewed area, the most widely used datasets are selected for benchmarking in Section V-A. Because of the high efficiency requirement for lane detection, speed tests are also conducted on representative open-source lane detection methods in a unified environment in Section V-B. Note that we only list published works for reference. Following the performance and efficiency comparisons, Section V-C revisits the existing methods according to the four core designs of lane detection.
@@ V-A Main Results on Lane Detection Datasets
CULane [36] and OpenLane [24] are currently the most widely used 2D and 3D lane detection datasets. We report the performance of representative methods on these two benchmarks in Table IV and Table V separately. All results are derived from the data in the original paper. More benchmark results are reported in Section -B.
@@ V-B Efficiency Comparison
Considering the variability of equipment used in experiments with existing methods, it is challenging to conduct a relatively fair comparison in terms of execution efficiency. Therefore, we retest representative methods in a unified environment. Table VI shows the work efficiency of these methods. The representative open-source methods are reevaluated according to their settings on the CULane or OpenLane dataset. To ensure fairness, only the inference speed of the model is tested to report the frames per second (FPS). The backbone, input size, model’s output, and possible post-processing (whether the model’s output reflects a vectorized representation of each unique lane instance) of each method are also described. All tests utilize a single Nvidia GeForce RTX 3090 GPU.
TABLE V: Benchmark results of 3D lane detection methods on OpenLane [24].
@@
TABLE VI: Efficiency comparison of representative methods. The setting of SCNN [36] and LSTR [82] is based on the re-implementation of Feng et al. [20]. 3D-LaneNet [22] is not open-source, so we test it according to Guo et al.’s reproduction [23]. Under Method, ‘$\dagger$’ denotes iterative regression of Anchor3DLane [25] and ”LATR-Lite” refers to the lite version of LATR [27].
@@
GANet [18] ResNet18 320$\times$800 106 Binary keypoint heatmap and offset map. Clustering
@@
CondLSTR [18] ResNet18 320$\times$800 47 Keypoint heatmap and offset map for each instance. Refinement
@@
BEV-LaneDet [26] ResNet34 576$\times$1024 83 BEV lane’s equidistant point coordinates and instance embedding, 3D lane heights. Clustering
@@ V-C Discussion
$\bullet$ Task Paradigm. segmentation-based methods achieve Instance-level discrimination and lane positioning in a two-stage approach. The majority of the algorithm’s runtime is occupied by independent instance discrimination processes. This makes them overall less efficient than object detection-based methods which are achieved in a one-stage. For object detection-based methods, it is worth considering the positive and negative sample matching strategy during network training, which will determine whether NMS is needed for post-processing.
$\bullet$ Lane modeling. In mask-based modeling methods [36, 56, 16], each pixel is classified, which can lead to inaccurate segmentation masks that subsequently hinder vectorized fitting. Thus, achieving optimal performance and efficiency remains challenging. In contrast, keypoints-based modeling, line anchor-based modeling, and curve-based modeling methods learn fewer points or parameters, directly yielding the vectorized results needed downstream.
@@
Curve-based modeling methods [82, 20] exhibit decent efficiency but fall short in terms of competitive performance on 2D lane detection benchmarks. Interestingly, this kind of method achieves strong results in 3D lane detection [130, 125]. As analyzed by Han et al. [140], this discrepancy is due to the ground height influence, which makes fitting irregular lanes challenging in the frontal view (FV). In contrast, these lanes appear smooth in the bird’s-eye view (BEV), where they can be more easily fitted.
@@
$\bullet$ Global context supplementing. Regardless of the genre, most methods converge on the consensus that supplementing global information significantly enhances lane detection performance, particularly for detecting obstructed lanes. Additionally, it is crucial to ensure that these specially designed structures achieve a balance between efficient processing and effective results. While this aspect has received limited attention in existing 3D lane detection benchmarks and methods, in practical applications, certain solutions in 2D lane detection can provide valuable references or be seamlessly integrated into 3D lane detection frameworks.
$\bullet$ Perspective effect elimination. The ultimate goal remains to achieve precise 3D lane detection to support downstream applications. Using IPM to project 2D lane detection results into 3D space is feasible. However, the assumption of a flat ground often yields incorrect results in BEV, even if predictions are accurate in FV. While projecting 2D lane detection into 3D space based on depth values [44] is straightforward, this approach depends heavily on depth estimation and cannot be optimized in an end-to-end manner.
Early 3D lane detection methods [22, 23], which still assume a flat ground, leverage IPM to construct BEV features. Some later approaches [24, 26, 125] improve on this by incorporating learnable methods, leading to enhanced performance. Alternatively, other methods [119, 25, 27] avoid BEV feature construction entirely, modeling 3D lanes directly and employing a 3D-to-2D forward projection to circumvent the inherent errors introduced by IPM. It should be pointed out that some methods employ Transformers to achieve view transformation of BEV features [24] or enable interaction between 3D lanes and FV features [27], resulting in improved performance. Nonetheless, the hardware deployment of advanced operators, such as deformable attention [99], remains a problem in need of optimization.
@@
Figure 8: Roadmap for the Development of Hotspots in Lane Detection Research. In each domain, a symbolic event is identified as a milestone and presented in chronological order. Each task includes popular datasets and methods to date.
@@ VI Extended Works of Lane Detection
There are also some works that have achieved rich results and received widespread attention in recent years, which are closely related to lane detection. In terms of task flow, they can be regarded as an upgrade in monocular image lane detection. We provide a brief introduction to them in this section. Figure 8 depicts the roadmap of the evolution from lane detection to its expansion works.
@@ VI-A Multi-task Perception
In autonomous driving, multiple perception tasks often need to be processed synchronously, in real-time, and parallel. A shared backbone can largely save computation costs and improve efficiency. Thus, leveraging a unified framework to conduct multiple perception tasks simultaneously gradually becomes a research hotspot. Early works [142, 143, 159, 160, 161, 162] connect multiple specific task heads after the feature extractor to simultaneously complete three tasks on BDD100K [38] dataset: object detection, drivable area segmentation, and lane detection. These methods achieve significant results in each task, which benefit from the powerful and efficient encoder and carefully designed multitasking learning strategy. However, the GT of lanes in the BDD100K are only annotated at the semantic-level, and usually only binary segmentation methods can be used. For lanes, further post-processing is needed to distinguish instances. Recent research mainly focuses on multi-task 3D perception using multiple cameras as sensors. PETRv2 [144] design detection query, segmentation query, and lane query to support 3D object detection, BEV segmentation, and 3D lane detection simultaneously. Li et al. [145] propose a unified representation method for multiple perception tasks. They represent 3D objects and 3D lanes as a kind of 3D vector field, which allows them to leverage a single-head unified model to achieve multi-task perception.
@@ VI-B Video Lane Detection
As mentioned in the previous chapters, current works attempt to supplement more global information to better detect lanes with unclear visual clues. However, these methods rely on detectors that use single images. In autonomous driving systems, video frames are captured continuously. Therefore, the correlation between frames can be used to more reliably detect obscure lanes in the current frame. For 2D lane detection, Zou et al. [163] and Zhang et al. [164] use recursive neural networks to fuse the features of the current frame with those of several past frames. Zhang et al. [41] aggregate the features of the current frame and multiple past frames based on Transformer. Tabelini et al. [165] extract lane features from video frames using LaneATT [80] and combine these features. Wang et al. [166] utilize spatiotemporal information from adjacent video frames by extending the feature aggregation module in RESA [16]. Jin et al. put RVLD [46], which includes an intra-frame lane detector to locate lanes in stationary frames and a predictive lane detector to use information from the previous frame for lane detection in the current frame. OMR [152], developed by the same team, employs vehicle masks occupying lanes to interact with historical frames, further improving the accuracy of lane detection in the current frame. For 3D lane detection, STLane3D [167] proposes a multi-frame pre-alignment layer under the BEV space, which uniformly projects features from different frames onto the same ROI region. Anchor3DLane-T [25] incorporates temporal information by projecting the 3D anchors of the current frame onto previous frames to sample features. CurveFormer++ [168] design a temporary Curve Cross Attention module based on CurveFormer [119], which can selectively utilize historical curve query and keypoints to propagate historical information frame by frame.
@@ VI-C Online HD Map Construction
HD maps are an essential module for autonomous driving. Although the traditional offline method of building HD maps can generate accurate map information and is adopted by many autonomous driving companies, it requires a lot of manual annotation costs. As an alternative, Some work try to design a novel HD map learning framework that makes use of on-car sensors and computation to estimate vectorized local semantic maps.
@@
HDMapNet [158] adopts a basic bottom-up segmentation approach to perform semantic segmentation on all map elements. Then it combines instance embedding and post-processing clustering to obtain each map element instance. This rasterization result still requires post-processing for downstream use, therefore, subsequent work attempts to predict vectorized maps end-to-end. BeMapNet [169] model map elements as segmented Bézier curves. It detects map elements first and then regresses the detailed points with a piecewise Bézier head. VectorMapNet [170] uses a line to represent all map elements and defines a hierarchical query representation. The points of the map elements are autoregressively output through the transformer decoder. However, it outputs the point set through autoregression, which leads to low efficiency. MapTR [148] design a unique representation method for map elements. It uses lines and polygons with uniform sampling points to represent line and area elements, respectively. Therefore, all map elements are represented as sets with the same number of points and different arrangement orders. Owing to its unified permutation-equivalent modeling approach and hierarchical query design, MapTR achieves advanced performance and efficiency on the nuScenes [146] dataset solely with camera input, providing a solid baseline for follow-up research. Afterward, MapTRv2 [171] improves self-attention and cross-attention in the decoder of MapTR, further enhancing both the accuracy and performance. PivotNet [149] proposes an end-to-end framework for representing map elements using pivot points. The purpose is to address the issue of shape information loss caused by using a fixed number and consistent position of points to represent complex map elements in MapTR. HIMap [150] meticulously designs feature extractors for MapTR’s hierarchical query, enabling the model to better learn instance-level features. StreamMapNet [172] improves MapTR in terms of timing. It overlays information from all historical frames together and implements a memory mechanism using recurrent late embedding. Then MapTracker [151] formalizes the online HD map construction as a tracking task and uses the history of memory latents to ensure consistency in reconstruction over time.

### v3

@@ IV-B BEV-based methods
where $F_{FV}$ denotes the FV feature. $F_{BEV}$ denotes the BEV feature which contains height information. $x$, $y$, $z$ denote coordinates in 3D space. $M_{trans}$ denotes view transformation module. $u$, $v$ denote corresponding pixel coordinates in terms of $x$, $y$, $z$. $[\bm{R\quad T}]$ and $\bm{K}$ are camera extrinsics and intrinsics.
@@ IV-B1 IPM for View Transformation
3D-LaneNet [22] is the first method which uses deep learning to predict 3D lanes directly from monocular images. The network first predicts the camera pitch angle and height to generate a differentiable IPM, combining the original FV feature map to create the BEV feature map. 3D-LaneNet demonstrates promising results in detecting 3D lanes from monocular images. Then Gen-LaneNet [23] directly uses 2D lane segmentation results as input for IPM, allowing for the utilization of extensive 2D lane data and enhancing the model’s generalization. In contrast to 3D-LaneNet, Gen-LaneNet offers more reliable supervision by using the camera pitch angle and height as GT. Li et al. [119] propose a new loss function based on Gen LaneNet to better extract the height information of 3D lanes from 2D lane representations. On the BEV plane, the above work uses vertical line anchors to model lanes. However, matching line anchors to ground truth is performed by measuring the distance at a predefined $y=Y_{ref}$, which may result in missed detections for short lanes. Therefore, 3D-LaneNet+ [130] avoids this issue using a bottom-up segmentation approach. Liu et al. [131] believe that the model can be independent of the ground truth camera pose provided by the benchmark. They design a two-stage network based on Transformer, which first predicts camera pose, i.e., the required parameters for IPM, then extracts BEV features, and finally regresses polynomial coefficients. 3D-SplineNet [123] treats lanes as B-spline curves on the BEV plane.
@@ IV-C1 Combining Depth Estimation
Similar to depth-assisted methods [135, 136, 137] in monocular 3D object detection, SALAD [44] decouples 3D lane detection into 2D lane segmentation and dense depth estimation tasks. With the help of estimated depth, the 2D lane coordinates can be projected into 3D space. Due to the availability of depth information, this method is independent of camera extrinsics.
TABLE IV: Benchmark results of the representative 2D lane detection methods on CULane [36]. For “Cross” scene, only false positives are shown. For convenient comparison, a boundary line is used to separate segmentation-based methods and object detection-based methods.
@@ IV-C2 Derictly Modeling 3D Lanes
The 3D object detection methods based on sparse BEV representations, such as DETR3D [138] and PETR [139], guide this approach. CurveFormer [120] constructs curve queries by modeling 3D lanes as 3D line anchors to provide explicit positional priors [140]. It designs a curve cross-attention mechanism to predict polynomial parameters for 3D lanes. Inspired by the line anchor feature pooling mechanism in LaneATT [81], Anchor3DLane [25] utilizes camera intrinsics and extrinsics to accurately project 3D line anchor points onto FV features. This facilitates anchor feature sampling, allowing the network to predict 3D coordinates and lane classification results based on the sampled anchor features. LATR [27] decomposes 3D line anchors into dynamically generated point-level and lane-level queries. It uses dynamic 3D ground position embeddings to interact with FV features, updating the lane query to bridge 3D space and 2D images. Dong et al. [125] model lanes as 3D Bézier curves and predict the curve control points via Transformer. Han et al. [141] express the 3D lane as a polynomial in 3D space. They design two Transformer structures to learn the 2D polynomial with the height of the X-O-Z plane and project the resulting 3D lane onto FV for supervised alignment. PVALane [124] generates sparser 3D line anchors than Anchor3DLane [25] by predicting 2D lanes in FV. It also introduces a module to align sampled FV and BEV features for more accurate 3D lane detection.
@@ V Benchmark Results
This section reports the performance of representative lane detection methods on commonly used public datasets. For each reviewed area, the most widely used datasets are selected for benchmarking in Section V-A. Because of the high efficiency requirement for lane detection, speed tests are also conducted on representative open-source lane detection methods in a unified environment which are shown in Section V-B. Note that we only list published works for reference. Following the performance and efficiency comparisons, Section V-C revisits the existing methods according to the four core designs of lane detection.
@@ V-A Main Results on Lane Detection Datasets
CULane [36] and OpenLane [24] are currently the most widely used 2D and 3D lane detection datasets. We report the performance of representative methods on these two benchmarks in Table IV and Table V separately. All results are derived from the data in the original paper. More benchmark results are reported in Section -B of Appendix.
@@ V-B Efficiency Comparison
Since the different methods are implemented on different platforms for the experiment, it is unfair to directly compare the speeds reported in their original papers. Therefore, we retest representative methods in a unified environment. Table VI shows the work efficiency of these methods. The representative open-source methods are reevaluated according to their settings on the CULane or OpenLane dataset. To ensure fairness, only the inference speed of the model is tested to report the frames per second (FPS). The backbone, input size, model’s output, and possible post-processing (whether the model’s output reflects a vectorized representation of each unique lane instance) of each method are also described. All tests are conducted on a single Nvidia GeForce RTX 3090 GPU.
TABLE V: Benchmark results of the representative 3D lane detection methods on OpenLane [24].
@@
TABLE VI: Efficiency comparison of representative methods. The setting of SCNN [36] and LSTR [83] is based on the re-implementation of Feng et al. [20]. 3D-LaneNet [22] is not open-source, so we tested it according to Guo et al.’s reproduction [23]. For Method, ‘$\dagger$’ denotes the iterative regression of Anchor3DLane [25] and ”LATR-Lite” refers to the lite version of LATR [27].
@@
GANet [18] ResNet18 320$\times$800 106 Keypoint heatmaps and offset maps of all lanes foreground. Clustering & Coordinates refinement
@@
CondLSTR [70] ResNet18 320$\times$800 47 Keypoint heatmaps and offset maps of each lane instance. Coordinates refinement
@@
BEV-LaneDet [26] ResNet34 576$\times$1024 83 BEV lane’s keypoint coordinates and instance embedding, 3D lane heights. Clustering
@@ V-C Discussion
$\bullet$ Task Paradigm. Segmentation-based methods achieve instance-level discrimination and lane positioning in a two-stage approach. The majority of the algorithm’s runtime is occupied by independent instance discrimination processes. This makes them overall less efficient than object detection-based methods which are achieved in one-stage. For object detection-based methods, it is necessary to consider the matching strategy of the positive and negative samples during the network training. This will determine whether NMS is needed for post-processing after the network inference.
$\bullet$ Lane modeling. In mask-based modeling methods [36, 56, 16], each pixel is classified, which can lead to inaccurate segmentation masks that subsequently hinder vectorized fitting. Thus, achieving optimal performance and efficiency remains challenging. In contrast, keypoints-based modeling, line anchor-based modeling, and curve-based modeling methods learn fewer points or parameters, directly yielding the vectorized results for downstream use.
@@
Curve-based modeling methods [83, 20] exhibit decent efficiency but fall short in terms of competitive performance on 2D lane detection benchmarks. Interestingly, this kind of method achieves strong results in 3D lane detection [131, 126]. As analyzed by Han et al. [141], this discrepancy is due to the ground height influence, which makes fitting irregular lanes challenging in FV. In contrast, these lanes appear smooth in BEV, where they can be more easily fitted.
@@
$\bullet$ Global context supplementing. Regardless of the genre, most methods converge on the consensus that supplementing global information significantly enhances lane detection performance, particularly for detecting occluded lanes. Additionally, it is crucial to ensure that these specially designed structures achieve a balance between efficient processing and effective results. While this aspect has received limited attention in existing 3D lane detection benchmarks and methods, in practical applications, certain solutions in 2D lane detection can provide valuable references or be seamlessly integrated into 3D lane detection frameworks.
$\bullet$ Perspective effect elimination. The ultimate goal persists in obtaining precise 3D lanes to support downstream applications. Using IPM to project 2D lane detection results into 3D space is feasible. However, the assumption of a flat ground often yields incorrect results in BEV, even if predictions are accurate in FV. While projecting the 2D lane detection results into 3D space based on depth values [44] is straightforward, this approach depends heavily on depth estimation and cannot be optimized in an end-to-end manner.
Early 3D lane detection methods [22, 23, 131], which still assume a flat ground, leverage IPM to construct BEV features. Some later approaches [24, 26, 126] improve on this by incorporating learnable ways, leading to enhanced performance. Alternatively, other methods [120, 25, 27] avoid BEV feature construction entirely, modeling 3D lanes directly and employing a 3D-to-2D forward projection to circumvent the inherent errors introduced by IPM. It should be pointed out that Transformer [142] has strong abilities in view transformation of BEV features [24] or interaction between 3D lanes and FV features [27]. This conclusion is also widely confirmed in related 3D object detection works [129, 139]. Nonetheless, the hardware deployment of advanced operators, such as deformable attention [100], remains a problem worth of optimization.
@@
Figure 8: The roadmap for the development of hotspots in lane detection research. In each domain, a symbolic event is identified as a milestone and presented in chronological order. The representative datasets and methods within each task are listed.
@@ VI Extended Works of Lane Detection
There are also some works that have received widespread attention in recent years, which are closely related to lane detection. In terms of task flow, they can be regarded as an upgrade on monocular image lane detection. We provide a brief introduction to them in this section. Figure 8 depicts a roadmap of the evolution from lane detection to its expansion works.
@@ VI-A Multi-task Perception
In autonomous driving, multiple perception tasks often need to be processed synchronously, in real-time, and in parallel. A shared backbone can save computation costs and improve efficiency greatly. Thus, leveraging a unified framework to conduct multiple perception tasks simultaneously gradually becomes a research hotspot. Early works [144, 145, 161, 162, 163, 164] connect multiple specific task heads after the feature extractor to simultaneously complete three tasks on the BDD100K [38] dataset: object detection, drivable area segmentation, and lane detection. These methods achieve impressive results in each task, which benefit from the powerful and efficient encoder and carefully designed multi-task learning strategy. However, the labels of lanes in BDD100K are only semantic-level annotations, and only binary segmentation methods can be used. Further post-processing is needed to distinguish each lane instance. Recent researches mainly focus on multi-task 3D perception. PETRv2 [146] designs detection query, segmentation query, and lane query to support 3D object detection, BEV segmentation, and 3D lane detection simultaneously. Li et al. [147] propose a unified representation method for multiple perception tasks. They represent 3D objects and 3D lanes as a kind of 3D vector field, which allows them to leverage a single-head unified model to achieve multi-task perception.
@@ VI-B Video Lane Detection
As mentioned in the previous chapters, current works attempt to supplement more global information to better detect lanes with unclear visual clues. However, these methods rely on detectors that use single images. In autonomous driving systems, video frames are captured continuously. Therefore, the correlation between frames can be used to more reliably detect obscure lanes in the current frame. For 2D lane detection, Zou et al. [165] and Zhang et al. [166] use recursive neural networks to fuse the features of the current frame with those of several past frames. Zhang et al. [41] aggregate the features of the current frame and multiple past frames based on Transformer. Tabelini et al. [167] extract lane features from video frames using LaneATT [81] and combine these features. Wang et al. [168] utilize spatiotemporal information from adjacent video frames by extending the feature aggregation module in RESA [16]. Jin et al. put RVLD [46], which includes an intra-frame lane detector to locate lanes in stationary frames and a predictive lane detector to use information from the previous frame for lane detection in the current frame. OMR [154] employs vehicle masks occupying lanes to interact with historical frames, further improving the accuracy of lane detection in the current frame. For 3D lane detection, STLane3D [169] proposes a multi-frame pre-alignment layer under the BEV space, which uniformly projects features from different frames onto the same ROI region. Anchor3DLane-T [25] incorporates temporal information by projecting the 3D anchors of the current frame onto previous frames to sample features. CurveFormer++ [170] designs a temporary Curve Cross Attention module based on CurveFormer [120], which can selectively utilize historical curve query and keypoints to propagate historical information frame by frame.
@@ VI-C Online HD Map Construction
HD maps are an essential module for autonomous driving. Although the traditional offline method of building HD maps can generate accurate map information and is adopted by many autonomous driving companies, it requires a lot of manual annotation costs. As an alternative, an increasing number of works try to design a novel HD map learning framework that makes use of on-car sensors and computation to estimate vectorized local semantic maps.
@@
HDMapNet [160] adopts a basic bottom-up segmentation approach to perform semantic segmentation on all map elements. Then it combines instance embedding and post-processing clustering to obtain each map element instance. This rasterization result still requires post-processing for downstream use, therefore, subsequent work attempts to predict vectorized maps end-to-end. BeMapNet [171] models map elements as segmented Bézier curves. It detects map elements first and then regresses the detailed points with a piecewise Bézier head. VectorMapNet [172] uses a line to represent all map elements and defines a hierarchical query representation. The points of the map elements are autoregressively output through the transformer decoder. However, it outputs the point set through autoregression, which leads to low efficiency. To solve this problem, MapTR [150] designs a unique representation method for map elements. It uses lines and polygons with uniform sampling points to represent line and area elements, respectively. Therefore, all map elements are represented as sets with the same number of points and different arrangement orders. Owing to its unified permutation-equivalent modeling approach and hierarchical query design, MapTR achieves advanced performance and efficiency on the nuScenes [148] dataset solely with camera input, providing a solid baseline for follow-up research. Afterward, MapTRv2 [173] improves self-attention and cross-attention in the decoder of MapTR, further enhancing both the accuracy and performance. PivotNet [151] proposes an end-to-end framework for representing map elements using pivot points. The purpose is to address the issue of shape information loss caused by using a fixed number and consistent position of points to represent complex map elements in MapTR. HIMap [152] meticulously designs feature extractors for MapTR’s hierarchical query, enabling the model to better learn instance-level features. StreamMapNet [174] improves MapTR in terms of timing. It overlays information from all historical frames together and implements a memory mechanism using recurrent late embedding. Then MapTracker [153] formalizes the online HD map construction as a tracking task and uses the history of memory latents to ensure consistency in reconstruction over time.